# TellWhisper: Tell Whisper Who Speaks When

### v1

@@ Abstract
Multi-speaker automatic speech recognition (MASR) aims to predict “who spoke when and what” from multi-speaker speech, a key technology for multi-party dialogue understanding. However, most existing approaches decouple temporal modeling and speaker modeling when addressing “when” and “who”: some inject speaker cues before encoding (e.g., speaker masking), which can cause irreversible information loss; others fuse identity by mixing speaker posteriors after encoding, which may entangle acoustic content with speaker identity. This separation is brittle under rapid turn-taking and overlapping speech, often leading to degraded performance. To address these limitations, we propose TellWhisper, a unified framework that jointly models speaker identity and temporal within the speech encoder. Specifically, we design TS-RoPE, a time-speaker rotary positional encoding: time coordinates are derived from frame indices, while speaker coordinates are derived from speaker activity and pause cues. By applying region-specific rotation angles, the model explicitly captures per-speaker continuity, speaker-turn transitions, and state dynamics, enabling the attention mechanism to simultaneously attend to “when” and “who”. Moreover, to estimate frame-level speaker activity, we develop Hyper-SD, which casts speaker classification in hyperbolic space to enhance inter-class separation and refine speaker-activity estimates. Extensive experiments demonstrate the effectiveness of the proposed approach.
@@
Yifan Hu1,2, Peiji Yang2, Zhisheng Wang2, Yicheng Zhong2, Rui Liu1thanks: Corresponding author. 1 Inner Mongolia University, Hohhot, China 2 Tencent Technology Co.Ltd, Shenzhen, China 22309013@mail.imu.edu.cn, imucslr@imu.edu.cn, {peijiyang, plorywang, ajaxzhong}@tencent.com
@@ 1 Introduction
Multi-speaker Automatic Speech Recognition (MASR) aims to predict who speaks what content and at what time in speech containing interactions among multiple speakers Polok et al. (2025); Yin et al. (2025). It is a complex task that jointly integrates speaker diarization (SD) Bredin et al. (2020) and automatic speech recognition (ASR) Cao et al. (2012). With the development of speech intelligence and conversational systems, MASR plays an increasingly critical role in meeting and interview transcription Vinnikov et al. (2024), multi-user human–computer interaction Shin et al. (2025), and the construction of data for spoken dialogue speech foundation models Ju et al. (2025); Xie et al. (2025). Consequently, developing efficient and robust MASR models is of practical importance.
@@
In summary, the main contributions of this paper are as follows: (1) We propose TellWhisper, a novel multi-speaker ASR model that introduces TS-RoPE, a time–speaker-aware rotary positional encoding, into the speech encoder to naturally integrate temporal and speaker activity. (2) To obtain reliable frame-level speaker activity, we develop Hyper-SD, a hyperbolic-space speaker diarization model that estimates speaker activity via “feature-prototype distances.” (3) We conduct extensive experiments that demonstrate the effectiveness of TS-RoPE for time-speaker integration and show that Hyper-SD provides reliable speaker-activity estimates.
@@ 6.1 Verifying the Reliability of Hyper-SD
6.2 Evaluating the Performance of Multi-Speaker Speech Recognition
@@ A.2 Experimental Setup
As shown in Table 8, we report the key hyperparameters of the main modules in TellWhisper. During training, we adopt different optimizers and learning rates for different components. (1) Speaker Activity Estimation (Hyper-SD). Optimizing parameters in hyperbolic space is a manifold-constrained problem with curvature, where standard Adam/AdamW (which performs Euclidean gradient updates) may lead to incorrect update directions, drifting off the manifold, and numerical instability. Therefore, for the hyperbolic speaker prototypes, we use Riemannian Adam, which performs Adam-style updates on the hyperbolic manifold, resulting in more stable optimization and faster convergence. The learning rate is set to $1\times 10^{-3}$. For the WavLM parameters, we use AdamW with a learning rate of $2\times 10^{-5}$; all remaining parameters are optimized with AdamW using a learning rate of $1\times 10^{-3}$. (2) Speaker-Time Aware Encoder and Structured Content Predictor. We use AdamW with a learning rate of $1\times 10^{-5}$ and $\epsilon=1\times 10^{-8}$.
@@ B.2.1 TellWhisper Performance on Overlapping Speech
As shown in Fig. 5 and  6, we conduct a qualitative case study on LibriCSS to examine model behavior under varying overlap ratios (0%–30%), focusing on speaker assignment, temporal alignment, and content transcription. Overall, TellWhisper remains robust as overlap increases and continues to produce coherent, well-structured outputs.
@@
In terms of content, the predicted transcripts largely preserve the semantics of the ground truth, with mismatches typically limited to occasional word-level substitutions in highly overlapped regions. Regarding temporal alignment, the model generally provides reasonable start/end boundaries. Higher overlap may lead to slightly finer-grained segmentation or minor boundary shifts, yet the overall timing remains well aligned. For speaker attribution, predictions are consistently accurate under low-to-moderate overlap, while the few confusions observed at higher overlap are mostly localized around overlap windows and do not substantially disrupt the global conversational structure. Taken together, these visualizations suggest that although heavy overlap increases local ambiguity, our TellWhisper maintains strong performance across speaker, time, and content dimensions, demonstrating good robustness under challenging multi-speaker conditions.

### v3

@@ Abstract
Multi-speaker automatic speech recognition (MASR) aims to predict “who spoke when and what” from multi-speaker speech, a key technology for multi-party dialogue understanding. However, most existing approaches decouple temporal modeling and speaker modeling when addressing “when” and “who”: some inject speaker cues before encoding (e.g., speaker masking), which can cause irreversible information loss; others fuse identity by mixing speaker posteriors after encoding, which may entangle acoustic content with speaker identity. This separation is brittle under rapid turn-taking and overlapping speech, often leading to degraded performance. To address these limitations, we propose TellWhisper, a unified framework that jointly models speaker identity and temporal within the speech encoder. Specifically, we design TS-RoPE, a time-speaker rotary positional encoding: time coordinates are derived from frame indices, while speaker coordinates are derived from speaker activity and pause cues. By applying region-specific rotation angles, the model explicitly captures per-speaker continuity, speaker-turn transitions, and state dynamics, enabling the attention mechanism to simultaneously attend to “when” and “who”. Moreover, to estimate frame-level speaker activity, we develop Hyper-SD, which casts speaker classification in hyperbolic space to enhance inter-class separation and refine speaker-activity estimates. Extensive experiments demonstrate the effectiveness of the proposed approach. The project webpage is available at https://walker-hyf.github.io/TellWhisper.
@@
Yifan Hu1,2, Peiji Yang2, Zhisheng Wang2, Yicheng Zhong2, Rui Liu1thanks: Corresponding author. 1 Inner Mongolia University, Hohhot, China 2 Tencent Technology Co.Ltd, Shenzhen, Guangdong, China 22309013@mail.imu.edu.cn, imucslr@imu.edu.cn, {peijiyang, plorywang, ajaxzhong}@tencent.com
@@ 1 Introduction
Multi-speaker Automatic Speech Recognition (MASR) aims to predict who speaks what content and at what time in speech containing interactions among multiple speakers Polok et al. (2025); Yin et al. (2025). It is a complex task that jointly integrates speaker diarization (SD) Bredin et al. (2020) and automatic speech recognition (ASR) Cao et al. (2012). With the development of speech intelligence and conversational systems, MASR plays an increasingly critical role in meeting and interview transcription Vinnikov et al. (2024), multi-user human–computer interaction Shin et al. (2025), and the construction of data for spoken dialogue speech foundation models Ju et al. (2025); Xie et al. (2025); Hao et al. (2025); Weihua et al. (2026). Consequently, developing efficient and robust MASR models is of practical importance.
@@
In summary, the main contributions of this paper are as follows:
  • We propose TellWhisper, a novel multi-speaker ASR model that introduces TS-RoPE, a time–speaker-aware rotary positional encoding, into the speech encoder to naturally integrate temporal and speaker activity.
  • To obtain reliable frame-level speaker activity, we develop Hyper-SD, a hyperbolic-space speaker diarization model that estimates speaker activity via “feature-prototype distances.”
  • We conduct extensive experiments that demonstrate the effectiveness of TS-RoPE for time-speaker integration and show that Hyper-SD provides reliable speaker-activity estimates.
@@ 6.1 Verifying the Reliability of Hyper-SD
6.2 Evaluating the Performance of Multi-Speaker Speech Recognition
@@ 7 Conclusion
Limitations
(1) The number of supported speakers is limited. In this work, TS-RoPE is designed for 1-4 speaker scenarios, consistent with the settings in SortFormer and DiCoW. We further extend the model to accommodate more speakers in future work. (2) The geometric consistency of Hyper-SD remains to be improved. At present, we perform hyperbolic classification only after feature extraction, leaving the encoder and the hyperbolic classifier in mismatched embedding spaces. In future work, we investigate end-to-end hyperbolic learning that encodes features directly in hyperbolic space using hyperbolic neural networks.
Ethics Statement
(1) About the datasets: All datasets used in this work are sourced from publicly available corpora and are accessed only after obtaining approval through the official application processes on their respective websites. We use only the audio recordings and the transcripts, and the data do not contain personally identifying information or offensive content. (2) About AI assistants: During model development, we use AI tools to assist with code writing, and all AI-generated code is subsequently reviewed and verified by humans. During paper writing, we use AI only for grammar checking.
Acknowledgment
This research of Rui Liu was funded by the General Program (No.62476146) of the National Natural Science Foundation of China, the Young Elite Scientists Sponsorship Program by CAST (2024QNRC001), the Outstanding Youth Project of Inner Mongolia Natural Science Foundation (2025JQ011), the Key R&D and Achievement Transformation Program of Inner Mongolia Autonomous Region (2025YFHH0014), the Central Government Fund for Promoting Local Scientific and Technological Development (2025ZY0143).
@@ A.2 Experimental Setup
As shown in Table 8, we report the key hyperparameters of the main modules in TellWhisper. During training, we adopt different optimizers and learning rates for different components. (1) Speaker Activity Estimation (Hyper-SD). Optimizing parameters in hyperbolic space is a manifold-constrained problem with curvature, where standard Adam/AdamW (which performs Euclidean gradient updates) may lead to incorrect update directions, drifting off the manifold, and numerical instability. Therefore, for the hyperbolic speaker prototypes, we use Riemannian Adam, which performs Adam-style updates on the hyperbolic manifold, resulting in more stable optimization and faster convergence. The learning rate is set to $1\times 10^{-3}$. For the WavLM parameters, we use AdamW with a learning rate of $2\times 10^{-5}$; all remaining parameters are optimized with AdamW using a learning rate of $1\times 10^{-3}$. (2) Speaker-Time Aware Encoder and Structured Content Predictor. We use AdamW with a learning rate of $1\times 10^{-5}$ and $\epsilon=1\times 10^{-8}$. All experiments are conducted on eight NVIDIA H20 GPUs.
@@ B.2.1 TellWhisper Performance on Overlapping Speech
As shown in Fig. 5 and  6, we conduct a qualitative case study on LibriCSS to examine model behavior under varying overlap ratios (0%–30%), focusing on speaker assignment, temporal alignment, and content transcription. Overall, TellWhisper remains robust as overlap increases and continues to produce coherent, well-structured outputs. In terms of content, the predicted transcripts largely preserve the semantics of the ground truth, with mismatches typically limited to occasional word-level substitutions in highly overlapped regions. Regarding temporal alignment, the model generally provides reasonable start/end boundaries. Higher overlap may lead to slightly finer-grained segmentation or minor boundary shifts, yet the overall timing remains well aligned. For speaker attribution, predictions are consistently accurate under low-to-moderate overlap, while the few confusions observed at higher overlap are mostly localized around overlap windows and do not substantially disrupt the global conversational structure. Taken together, these visualizations suggest that although heavy overlap increases local ambiguity, our TellWhisper maintains strong performance across speaker, time, and content dimensions, demonstrating good robustness under challenging multi-speaker conditions.